Suppose, a lighthouse operator (L) observes a shipping boat (B) that is moving straight away from the light house. The level of the lighthouse operator is 20 m above the level of the sea. At a particular instant, the lighthouse operator observe the boat (B1) with a depression angle of 400. Two minutes later, the boat (B2) was observed with a depression angle of 800.
• What is vB, the speed of the boat as measured by the light house operator in m/s?
(You can use speed = distance relation to determine the speed.) time
2
1
Expert's answer
2019-09-17T12:09:06-0400
"d_1=h\tan{40\degree}"

"d_2=h\tan{80\degree}"

The speed of the boat as measured by the light house operator:


"v_B=\frac{d_2-d_1}{t}=\frac{h(\tan{80\degree}-\tan{40\degree})}{t}"

"v_B=\frac{20(\tan{80\degree}-\tan{40\degree})}{2(60)}=0.8\frac{m}{s}"